OpenStudy (anonymous):

what is the inverse function of x^3+5 ??

OpenStudy (2bornot2b):

\[\sqrt[3]{x-5}\]

OpenStudy (anonymous):

y = x^3 + 5 Finding the inverse, we make x as y and y as x x = y^3 -5 And we isolate y, by adding 5 to each side x+5 = y^3 Then taking the cube root of each side, getting what 2bornot2b got